can u solve this equation for me ?
2^2x - 11.2^3 + 15 = 0
im stucked on it please help me thanks in advance ;)

Answer by Alan3354(69443)

2^2x - 11.2^3 + 15 = 0
2^2x = 11.2^3 - 15 = 1389.928
2x*log(2) = log(1389.28)
2x = log(1389.28)/log(2)
2x =~ 10.4408
x =~ 5.2204
-----------
I doubt that this was the problem you meant to post.